Two bond funds pay interest at rates that differ by 2%. Money invested for one year in the first fund earns $225 interest. The same amount invested in the second fund earns $275. Find the lower rate of interest.
Answer by jorel1380(3719)

Let n be the amount invested in each of these two bond fund, and r be the lower interest rate. Then the higher interest rate would be r+.02. So:
n(r)=225
n(r+.02)=275
So:
225/r=275/r+.02
275r=225r+4.5
50r=4.5
r=4.5/50=.09
The lower interest rate is .09, or 9%;while the higher interest rate is 11%. ☺☺☺☺